OpenStudy (anonymous):

With the purchase of 10 greeting cards, Addie recieved a discount. She paid $26.35 for the cards that would normally have cost $31. What percent discount did Addie recieve?

OpenStudy (anonymous):

15% $31/10 cards=3.1 $26.35/10 cards=2.635 2.635/3.1=.85 1-.85=.15